OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 3 - 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (ajprincess):

Do you knw distributive property?

OpenStudy (ajprincess):

@courtneyaslan

OpenStudy (anonymous):

yeah I know how to do it i have to 1.distribute 2.do like terms 3.keep variable on one side, number on other 4.solve for x 5.check by plugging answer onto original problem but i cant solve the answer

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

why did u call me dear? lol

OpenStudy (anonymous):

oh lol so as i said i know how to find the answer but i cant can u help me @ajprincess

OpenStudy (ajprincess):

1.distribute 3 - 5(x + 2) = 3(x + 1) + 6 3-5*x-5*2=3*x+3*1+6 2.do like terms 3-5x-10=3x+3+6 3.keep variable on one side, number on other -5x-7=3x+9 -7-9=3x+5x 8x=-16 4.solve for x x=-16/8 x=-2 5.check by plugging answer onto original problem RHS=> 3 - 5(x + 2) =3-5(-2+2)=3-5*0=3 LHS=> 3(x + 1) + 6 =3(-2+1)+6=3(-1)+6=-3+6=3 RHS=LHS so x=-2 Nw what ptoblem do u hav in this.

OpenStudy (anonymous):

omg thx no i see the mistake i made thankyou so much

OpenStudy (ajprincess):

yw.

OpenStudy (anonymous):

:D :P